OpenStudy (anonymous):

lim as x approaches infinity of xtan(3/x)

OpenStudy (anonymous):

God these are giving me so much trouble... okay here's my attempt.

OpenStudy (anonymous):

Should I turn it into a fraction before I do lhopitals? As in, turn tan in to sin/cos???

myininaya (myininaya):

\[\text{ Let } u=\frac{3}{x} =>\text{ as } x->\infty, u-> 0\] => x=3/u \[\lim_{u \rightarrow 0}\frac{3}{u} \cdot \tan(u)\] we can apply l'hosptial rule since we have 0/0 \[3 \lim_{u \rightarrow 0}\frac{\sec^2(u)}{1}=3 \sec^2(0)=3(1)=3\]

OpenStudy (lalaly):

let \[u=\frac{1}{x} \] \[\lim_{u \rightarrow 0} \frac{\tan3u}{u} \]using lhopitals rule \[\lim_{u \rightarrow 0}\frac{3\sec^2u}{1}\]\[=\lim_{u \rightarrow 0}\frac{3}{\cos^2{3u}} =\frac{3}{1} = 3\]

OpenStudy (anonymous):

I suppose I don't need to. Lhopitals should work regardless of if it's in fraction form. Okay so.... xtan(3/x) = infinity * 0 when x is approaching infinity. That's a prerequisite of lhopitals! SO i can get right down to business. So derivative of x is 1. Then derivative of tan(3/x) is sec^2(3/x) *(-3/x^2) Did i do that correctly ??^^^

OpenStudy (anonymous):

@myininaya I dont get how the hell you got (3/u).... from the (3/u)tan(u)

myininaya (myininaya):

i let u=3/x so ux=3 so x=3/u if u=3/x and x goes to infinity, then u goes to zero